OpenStudy (anonymous):

Solve log7 (x – 8) – log7 3 = log7 21

OpenStudy (anonymous):

Helllllpppppp

OpenStudy (anonymous):

log7(x-8)-log7(3) = log7(21) log7[(x-8)/(3)] = log7(21) (x-8)/3=21 x-8=63 x=71

OpenStudy (anonymous):

When two logs, which have the same base, are being subtracted, we can divide them. \[\log_{7} (x-8) / 3 = \log_{7}21 \] Bring the log on the other side to the left so that zero is left on the right. And since the log will be subtracted, we will again divide the two.\[\log_{7} (x-8) / 63 = 0\]Now using the base to the power rule of logarithm, 7 will be raised to the power 0 which will be 1 and the inside of the log (x-8)/63 will appear without the log. \[1 = x-8\div63\] \[63 = x -8\]\[71 = x\]